OpenStudy (hennawho):

MEDAL!!! xtanx - 3tanx = x-3

OpenStudy (google):

oh i thought it said xanax

OpenStudy (faiqraees):

factor tan x on left side

OpenStudy (agent0smith):

\[\large x \tan x - 3 \tan x = (x-3)\]after factoring\[\large \tan x(x - 3) = (x-3)\]subtract x-3 on both sides, but keep it in parentheses\[\large \tan x(x - 3) - (x-3) = 0 \]notice x-3 is a common factor\[\large (x-3)(\tan x - 1) = 0 \]Now solve by setting each factor equal to zero.